# Hypergraph Foundation Model

### v1

@@ IV-B2 Hierachical Neighborhood Representation
${\bm{Z}}_{x}=\frac{{\bm{v}}_{x}}{||{\bm{v}}_{x}||},\text{ where }{\bm{v}}_{x}=\sum_{u\in{\mathcal{N}}_{v}(x)}\psi(T_{u}),\forall x\in[L],$ (6)
where $\psi$ represents the basic text feature extraction network, such as TF-IDF[45], and $T_{u}$ is the textual description of vertex $u$. The set ${\mathcal{N}}_{v}(x)$ comprises vertices associated with label $x$. The matrix ${\bm{Z}}\in{\mathbb{R}}^{L\times C}$ serves as the initial representation matrix for the labels.
@@
${\bm{P}}^{(t)}\leftarrow\text{k-means-clustering}({\bm{Z}}^{(t)},K^{(t-1)}),$ (7)
where $\mathbf{P}^{(t)}\in\{0,1\}^{K^{(t)}\times K^{(t-1)}}$ is the projection matrix mapping clusters from layer $t$ to layer $t-1$, and $K^{(t)}$ denotes the number of clusters at layer $t$. The hierarchical process is repeated until the desired tree height $T$ is achieved, with the finest granularity at layer $T$ where $K^{(T)}=L$ and ${\bm{Z}}^{(T)}={\bm{Z}}$. For coarser layers, the label representations are updated as follows:
${\bm{Z}}^{(t-1)}={\bm{Z}}^{(t)}\mathbf{P}^{(t)}$ (8)
where ${\bm{Z}}^{(t-1)}\in{\mathbb{R}}^{K^{(t-1)}\times C}$ represents the aggregated label features at the $(t-1)$-th layer. Each row in ${\bm{Z}}^{(t-1)}$ corresponds to a group of vertices from the previous layer, effectively capturing higher-order neighborhood structures.
@@
${\bm{Y}}^{(t-1)}=\text{binary}\left({\bm{Y}}^{(t)}\mathbf{P}^{(t)}\right),$ (10)
where ${\bm{Y}}^{(t)}\in\{0,1\}^{N\times K^{(t)}}$, and the binary function binarizes the resulting matrix by setting all non-zero entries to 1. This hierarchical approach aggregates similar neighborhoods, creating a multi-level neighborhood representation. Consequently, structural information is progressively injected into the LM during training, enabling the handling of larger-scale hypergraphs and improving both encoding and computational efficiency.
@@ IV-B3 Association Information Injection
Specifically, we begin at the first level and iteratively predict each vertex’s associations with other vertices or vertex clusterings. At each level $t$, we first obtain the predicted associations for the $(t-1)$-th level, denoted as ${\bm{Y}}^{(t-1)}_{\text{pred}}$. To enhance training efficiency and robustness, inspired by beam search, we select the top $k$ most significant association labels for each vertex:
${\bm{Y}}^{(t-1)}_{\text{pred}}=\text{Top}\left({\bm{W}}^{(t-1)\top}\Phi({\mathcal{T}},\Theta),\ k\right)$ (11)
where ${\bm{Y}}^{(t-1)}_{\text{pred}}\in\{0,1\}^{N\times K^{(t-1)}}$ indicates whether each vertex is associated with the top $k$ labels at level $t-1$. However, utilizing the entire vector for training would lead to an imbalance between positive and negative samples, resulting in suboptimal model performance. To mitigate this issue, we introduce a mask that selects a subset of samples for training at each level:
${\bm{S}}^{(t)}=\text{binarize}({\bm{Y}}^{(t-1)}_{\text{pred}}{\bm{P}}^{(t)\top}+{\bm{Y}}^{(t-1)}{\bm{P}}^{(t)\top})$ (12)
where ${\bm{S}}^{(t)}\in\{0,1\}^{N\times K^{(t)}}$ is the sample selection matrix. The first term, ${\bm{Y}}^{(t-1)}_{\text{pred}}{\bm{P}}^{(t)\top}$, represents the associations predicted from the previous level, while the second term, ${\bm{Y}}^{(t)}{\bm{P}}^{(t)\top}$, denotes the true associations at the current level. The binarize function converts the resulting matrix into a binary form, setting all non-zero entries to 1 and others to 0.
@@
$\underset{{\bm{W}}^{(t)},\Theta}{\min}\sum\limits_{i=1}^{N}\sum\limits_{x:{\bm{S}}^{(t)}_{i,x}\neq 0}{\mathcal{L}}_{\text{BCE}}\left({\bm{Y}}^{(t)}_{i,x},{\bm{W}}^{(t)\top}_{x}\Phi({\mathcal{T}},\Theta)\right)$ (13)
@@ IV-C Hierarchical Multi-Hypergraph Guided Structural Knowledge Extraction
Figure 4: Illustration of sampling and building Hierarchical Multi-Hypergraphs for the pertaining of hypergraph foundation model.
@@ IV-C1 Multi-Domain Hypergraphs Sampling
[…]
@@
${\mathcal{G}}^{a}_{\text{sub}}$ =$=\{{\mathcal{V}}^{a}_{\text{sub}},{\mathcal{E}}^{a}_{\text{sub}}\}\quad\text{where}$ (16)
@@ IV-C2 Multi-Domain Hypergraphs Clustering
$\{c_{1}^{a},c_{2}^{a},\cdots,c_{k}^{a}\}\leftarrow\text{k-means-clustering}({\bm{X}}_{\text{sub}}^{a},k),$ (17)
@@
${\mathcal{V}}^{a}_{\text{clu}}&=\{v^{\text{clu}}_{a,1},v^{\text{clu}}_{a,2},\cdots,v^{\text{clu}}_{a,k}\},\\ {\mathcal{E}}^{a}_{\text{clu}}&=\text{knn}({\bm{X}}^{a}_{\text{clu}},k_{c})\cup\{c^{a}_{1},\cdots,c^{a}_{k}\}\end{aligned}\right.,$ (19)
@@
IV-C3 Hierachical Multi-Hypergraph Construction
@@
${\mathcal{V}}_{\text{bond}}&=\{v^{\text{bond}}_{a},v^{\text{bond}}_{b},\cdots,v^{\text{bond}}_{m}\}\\ {\mathcal{E}}_{\text{bond}}&=\{(v^{\text{bond}}_{a},v^{\text{bond}}_{b},\cdots,v^{\text{bond}}_{m}),\\ &\qquad(v^{\text{bond}}_{a},v^{\text{clu}}_{a,1},v^{\text{clu}}_{a,2},\cdots,v^{\text{clu}}_{a,k}),\\ &\qquad(v^{\text{bond}}_{b},v^{\text{clu}}_{b,1},v^{\text{clu}}_{b,2},\cdots,v^{\text{clu}}_{b,k}),\\ &\qquad\qquad\cdots\\ &\qquad(v^{\text{bond}}_{m},v^{\text{clu}}_{m,1},v^{\text{clu}}_{m,2},\cdots,v^{\text{clu}}_{m,k})\}\end{aligned}\right.$ (20)
@@
${\bm{x}}^{\text{bond}}_{d}=\frac{1}{k}\sum_{j=1}^{k}{\bm{x}}^{\text{clu}}_{d,c_{j}}$ (21)
@@
${\bm{X}}_{\text{bond}}={\bm{x}}^{\text{bond}}_{a}||{\bm{x}}^{\text{bond}}_{b}||\cdots||{\bm{x}}^{\text{bond}}_{m}$ (22)
@@
${\mathcal{G}}_{\text{M}}&=\{{\mathcal{V}}_{\text{M}},{\mathcal{E}}_{\text{M}}\},\\ {\mathcal{V}}_{\text{M}}&={\mathcal{V}}_{\text{bond}}\cup\bigcup_{d=1}^{m}\left({\mathcal{V}}_{\text{sub}}^{d}\cup{\mathcal{V}}_{\text{clu}}^{d}\right),\\ {\mathcal{E}}_{\text{M}}&={\mathcal{E}}_{\text{bond}}\cup\bigcup_{d=1}^{m}\left({\mathcal{E}}_{\text{sub}}^{d}\cup{\mathcal{E}}_{\text{clu}}^{d}\right),\\ {\bm{X}}_{\text{M}}&={\bm{X}}_{\text{bond}}||{\bm{X}}_{\text{sub}}^{a}||{\bm{X}}_{\text{clu}}^{a}||\cdots||{\bm{X}}_{\text{sub}}^{m}||{\bm{X}}_{\text{clu}}^{m}\end{aligned}\right.$ (23)
@@
This hierarchical multi-domain hypergraph structure is derived through the sampling of multiple domain-specific hypergraphs, thereby supporting the training of a multi-domain foundation model. During training, sampling can be performed in each epoch to ensure comprehensive coverage of the hypergraph information. Notably, the virtual vertices within the hierarchical multi-domain hypergraph—comprising both cluster vertices and bond vertices—are characterized by aggregated features rather than trainable embeddings. This design choice enhances the transferability of the model, enabling the construction of arbitrary multi-domain hypergraphs and supporting effective learning across diverse domains.
@@ IV-D Pretraining on Multi-Domain Hypergraphs
6:     $\{{\bm{P}}^{(t)}\}_{t=1}^{T}$leftarrow k-means-clustering$({\bm{Z}},{\mathcal{K}})$
7:     $\{{\bm{Y}}^{(t)}\}_{t=1}^{T}$leftarrow $\text{binary}({\bm{Y}}^{(t)}{\bm{P}}^{(t)})$
@@
21:     $\{({\mathcal{G}}^{j}_{\text{sub}},{\bm{X}}^{j}_{\text{sub}})\}_{j=1}^{m}$leftarrow sampling($(\{({\mathcal{G}}^{j},{\bm{X}}_{j})\}_{j=1}^{m})$
@@
24:         $\{c_{1}^{j},c_{2}^{j},\cdots,c_{k}^{j}\}$leftarrow k-means-clustering$({\bm{X}}^{j}_{\text{sub}},k)$
25:         $({\mathcal{V}}^{j}_{\text{clu}},{\mathcal{E}}^{j}_{\text{clu}})$leftarrow by Eq. 19
@@
where, the structural self-supervised loss ${\mathcal{L}}_{\text{stru}}$ is implemented using contrastive learning techniques applied to augmented structural views of the hypergraph, such as those employed in HyperGCL [18]. This encourages the model to learn robust structural representations by distinguishing between different structural augmentations. The feature self-supervised loss ${\mathcal{L}}_{\text{feat}}$ involves masking portions of the vertex features and reconstructing them based on domain-specific information, similar to the approach used in GraphMAE [46]. This mechanism ensures that the model effectively captures both the structural dependencies and the semantic features of the vertices.
@@ IV-F Discussions
In this subsection, we compare the proposed Hyper-FM with the classical graph foundation model Graph COordinators for PrEtraining (GCOPE)[47]. The comparison focuses on three key aspects: vertex feature generation, multi-domain association structure generation, and scalability and transferability. [48] for dimensionality reduction. Firstly, regarding vertex feature generation, GCOPE directly employs BoW [16] or TF-IDF [45] to generate vertex features, followed by SVD [17] to unify the feature dimensions. This approach does not fully leverage the capabilities of advanced LM for extracting textual features, potentially leading to performance limitations. In contrast, Hyper-FM utilizes LM to extract more effective and nuanced textual features, enhancing the quality of vertex representations. Furthermore, GCOPE’s feature extraction process does not incorporate the inherent relationships present in the original data. Hyper-FM addresses this limitation by leveraging hierarchical domain representations and incorporating a domain prediction task to fine-tune the LM. This enables Hyper-FM to extract textual features that are not only rich in semantic information but also cognizant of domain-specific encodings, thereby making more comprehensive use of the original hypergraph’s structural information.
Secondly, in the generation of multi-domain association structures, GCOPE concatenates multi-domain structural data using virtual vertices. These virtual vertices are connected to all vertices within a particular domain, which can lead to the blending and confusion of information among semantically distinct vertices within the same domain. Hyper-FM, on the other hand, adopts a clustering-based approach to construct hierarchical multi-domain hypergraph structures. This method enhances the internal connections among semantically similar vertices within each domain, effectively preventing the mixing of disparate semantic knowledge. Additionally, rather than connecting domain vertices directly to all constituent vertices, Hyper-FM connects them to clustered vertices. This strategy promotes the orderly transfer of knowledge across different domains, ensuring that the information flow is both structured and semantically meaningful.
Lastly, concerning scalability and transferability, GCOPE’s method of directly concatenating diverse domain graph structures poses significant challenges when dealing with large-scale hypergraphs. Hyper-FM mitigates this issue by sampling substructures from different domains in each training epoch. This not only captures the diverse structural associations inherent in multi-domain data but also equips the foundation model with the capacity to handle and learn from large-scale hypergraph structures effectively. Additionally, GCOPE initializes virtual vertex embeddings with trainable parameters tailored to each domain, necessitating complete retraining when new domains are introduced. In contrast, Hyper-FM generates virtual cluster vertex and bond vertex features through feature aggregation, eliminating the need for additional trainable parameters. This design choice enhances the model’s adaptability, allowing it to seamlessly incorporate new domains without necessitating extensive retraining. Consequently, Hyper-FM demonstrates superior scalability and transferability, making it well-suited for dynamic and evolving multi-domain hypergraph scenarios.
Overall, Hyper-FM presents significant advancements over GCOPE in effectively generating vertex features, constructing robust multi-domain association structures, and ensuring scalability and transferability. These enhancements underscore the efficacy of Hyper-FM in capturing and leveraging complex multi-domain hypergraph information, thereby setting a higher performance benchmark for hypergraph foundation models.

### v2

@@ IV-B2 Hierarchical Neighborhood Representation
[…]
where $\psi$ represents the basic text feature extraction network, such as TF-IDF[50]. It should be noted that $\psi$ only extracts initial textual features, which are not directly fed into downstream neural networks but rather serve as the basis for clustering to provide references for hierarchical domains. $T_{u}$ is the textual description of vertex $u$. The set ${\mathcal{N}}_{v}(x)$ comprises vertices associated with label $x$. The matrix ${\bm{Z}}\in{\mathbb{R}}^{L\times C}$ serves as the initial representation matrix for the labels.
@@
${\bm{P}}^{(t)}\leftarrow k\text{-means-clustering}({\bm{Z}}^{(t)},K^{(t-1)}){\color[rgb]{0,0,0},}$ (7)
where $\mathbf{P}^{(t)}\in\{0,1\}^{K^{(t)}\times K^{(t-1)}}$ is the projection matrix mapping clusters from layer $t$ to layer t1t-1, and $K^{(t)}$ denotes the number of clusters at layer $t$. The hierarchical process is repeated until the desired tree height $T$ is achieved, with the finest granularity at layer $T$ where $K^{(T)}=L$ and ${\bm{Z}}^{(T)}={\bm{Z}}$. For coarser layers, the label representations are updated as follows:
${\bm{Z}}^{(t-1)}={\bm{Z}}^{(t)}\mathbf{P}^{(t)}{\color[rgb]{0,0,0},}$ (8)
where ${\bm{Z}}^{(t-1)}\in{\mathbb{R}}^{K^{(t-1)}\times C}$ represents the aggregated label features at the (t1)(t-1)-th layer. Each row in ${\bm{Z}}^{(t-1)}$ corresponds to a group of vertices from the previous layer, effectively capturing higher-order neighborhood structures.
@@
${\bm{Y}}^{(t-1)}=\text{bin}\left({\bm{Y}}^{(t)}\mathbf{P}^{(t)}\right),$ (10)
where ${\bm{Y}}^{(t)}\in\{0,1\}^{N\times K^{(t)}}$, and the bin function binarizes the resulting matrix by setting all non-zero entries to 1. This hierarchical approach aggregates similar neighborhoods, creating a multi-level neighborhood representation. Consequently, structural information is progressively injected into the LM during training, enabling the handling of larger-scale hypergraphs and improving both encoding and computational efficiency.
@@ IV-B3 Association Information Injection
Specifically, we begin at the first level and iteratively predict each vertex’s associations with other vertices or vertex clusterings. At each level $t$, we first obtain the predicted associations for the (t1)(t-1)-th level, denoted as ${\bm{Y}}^{(t-1)}_{\text{pred}}$. To enhance training efficiency and robustness, inspired by beam search, we select the top $k$ most significant association labels for each vertex:
${\bm{Y}}^{(t-1)}_{\text{pred}}=\text{Top}\left({\bm{W}}^{(t-1)\top}\Phi({\mathcal{T}},\Theta),\ k\right){\color[rgb]{0,0,0},}$ (11)
where ${\bm{Y}}^{(t-1)}_{\text{pred}}\in\{0,1\}^{N\times K^{(t-1)}}$ indicates whether each vertex is associated with the top $k$ labels at level t1t-1. However, utilizing the entire vector for training would lead to an imbalance between positive and negative samples, resulting in suboptimal model performance. To mitigate this issue, we introduce a mask that selects a subset of samples for training at each level:
${\bm{S}}^{(t)}=\text{bin}({\bm{Y}}^{(t-1)}_{\text{pred}}{\bm{P}}^{(t)\top}+{\bm{Y}}^{(t-1)}{\bm{P}}^{(t)\top}){\color[rgb]{0,0,0},}$ (12)
where ${\bm{S}}^{(t)}\in\{0,1\}^{N\times K^{(t)}}$ is the sample selection matrix. The first term, ${\bm{Y}}^{(t-1)}_{\text{pred}}{\bm{P}}^{(t)\top}$, represents the associations predicted from the previous level, while the second term, ${\bm{Y}}^{(t)}{\bm{P}}^{(t)\top}$, denotes the true associations at the current level. The bin function converts the resulting matrix into a binary form, setting all non-zero entries to 1 and others to 0.
@@
$\underset{{\bm{W}}^{(t)},\Theta}{\min}\sum\limits_{i=1}^{N}\sum\limits_{x:{\bm{S}}^{(t)}_{i,x}\neq 0}{\mathcal{L}}_{\text{BCE}}\left({\bm{Y}}^{(t)}_{i,x},{\bm{W}}^{(t)\top}_{x}\Phi({\mathcal{T}},\Theta)\right){\color[rgb]{0,0,0},}$ (13)
@@ IV-C Hierarchical Multi-Hypergraph Guided Structural Knowledge Extraction
Figure 4: Illustration of sampling and building Hierarchical Multi-Hypergraphs for the pretaining
of hypergraph foundation model.
@@ IV-C1 Multi-Domain Hypergraphs Sampling
$\{{\mathcal{G}}^{1}_{\text{sub}},{\mathcal{G}}^{2}_{\text{sub}},\ldots,{\mathcal{G}}^{m}_{\text{sub}}\}\xleftarrow{\text{sampling}}\{{\mathcal{G}}^{1},{\mathcal{G}}^{2},\ldots,{\mathcal{G}}^{m}\}.$ (15)
@@
${\mathcal{G}}^{a}_{\text{sub}}$ $=\{{\mathcal{V}}^{a}_{\text{sub}},{\mathcal{E}}^{a}_{\text{sub}}\}\quad\text{where}$ (16)
@@ IV-C2 Multi-Domain Hypergraphs Clustering
$\{c_{1}^{a},c_{2}^{a},\cdots,c_{k}^{a}\}\leftarrow{\color[rgb]{0,0,0}k\text{-means-clustering}}({\bm{X}}_{\text{sub}}^{a},k),$ (17)
@@
$\left\{\begin{aligned} {\mathcal{V}}^{a}_{\text{clu}}&=\{v^{\text{clu}}_{a,1},v^{\text{clu}}_{a,2},\cdots,v^{\text{clu}}_{a,k}\},\\ {\mathcal{E}}^{a}_{\text{clu}}&=\text{knn}({\bm{X}}^{a}_{\text{clu}},k_{c})\cup\{c^{a}_{1},\cdots,c^{a}_{k}\}\end{aligned}\right$., (19)
@@
IV-C3 Hierarchical Multi-Hypergraph Construction
@@
$\left\{\begin{aligned} {\mathcal{V}}_{\text{bond}}&=\{v^{\text{bond}}_{a},v^{\text{bond}}_{b},\cdots,v^{\text{bond}}_{m}\}\\ {\mathcal{E}}_{\text{bond}}&=\{(v^{\text{bond}}_{a},v^{\text{bond}}_{b},\cdots,v^{\text{bond}}_{m}),\\ &\qquad(v^{\text{bond}}_{a},v^{\text{clu}}_{a,1},v^{\text{clu}}_{a,2},\cdots,v^{\text{clu}}_{a,k}),\\ &\qquad(v^{\text{bond}}_{b},v^{\text{clu}}_{b,1},v^{\text{clu}}_{b,2},\cdots,v^{\text{clu}}_{b,k}),\\ &\qquad\qquad\cdots\\ &\qquad(v^{\text{bond}}_{m},v^{\text{clu}}_{m,1},v^{\text{clu}}_{m,2},\cdots,v^{\text{clu}}_{m,k})\}\end{aligned}\right$. (20)
@@
${\bm{x}}^{\text{bond}}_{d}=\frac{1}{k}\sum_{j=1}^{k}{\bm{x}}^{\text{clu}}_{d,c_{j}}{\color[rgb]{0,0,0},}$ (21)
@@
${\bm{X}}_{\text{bond}}={\bm{x}}^{\text{bond}}_{a}||{\bm{x}}^{\text{bond}}_{b}||\cdots||{\bm{x}}^{\text{bond}}_{m}{\color[rgb]{0,0,0}.}$ (22)
@@
$\left\{\begin{aligned} {\mathcal{G}}_{\text{M}}&=\{{\mathcal{V}}_{\text{M}},{\mathcal{E}}_{\text{M}}\},\\ {\mathcal{V}}_{\text{M}}&={\mathcal{V}}_{\text{bond}}\cup\bigcup_{d=1}^{m}\left({\mathcal{V}}_{\text{sub}}^{d}\cup{\mathcal{V}}_{\text{clu}}^{d}\right),\\ {\mathcal{E}}_{\text{M}}&={\mathcal{E}}_{\text{bond}}\cup\bigcup_{d=1}^{m}\left({\mathcal{E}}_{\text{sub}}^{d}\cup{\mathcal{E}}_{\text{clu}}^{d}\right),\\ {\bm{X}}_{\text{M}}&={\bm{X}}_{\text{bond}}||{\bm{X}}_{\text{sub}}^{a}||{\bm{X}}_{\text{clu}}^{a}||\cdots||{\bm{X}}_{\text{sub}}^{m}||{\bm{X}}_{\text{clu}}^{m}\end{aligned}\right$. (23)
@@
This hierarchical multi-domain hypergraph structure is derived through the sampling of multiple domain-specific hypergraphs, thereby supporting the training of a multi-domain foundation model. During training, sampling can be performed in each epoch to ensure comprehensive coverage of the hypergraph information. Notably, the virtual vertices within the hierarchical multi-domain hypergraph, comprising both cluster vertices and bond vertices, are characterized by aggregated features rather than trainable embeddings. This design choice enhances the transferability of the model, enabling the construction of arbitrary multi-domain hypergraphs and supporting effective learning across diverse domains.
@@ IV-D Pretraining on Multi-Domain Hypergraphs
[…]
7:  $\{{\bm{Y}}^{(t)}\}_{t=1}^{T}$ $\leftarrow$ $\text{binary}({\bm{Y}}^{(t)}{\bm{P}}^{(t)})$
@@
[…]
@@
[…]
25:   $({\mathcal{V}}^{j}_{\text{clu}},{\mathcal{E}}^{j}_{\text{clu}})$ $\leftarrow$ by Eq. 19
@@
where, the structural self-supervised loss ${\mathcal{L}}_{\text{stru}}$ is implemented using contrastive learning techniques applied to augmented structural views of the hypergraph, such as those employed in HyperGCL [18]. Following the design in SS-HT[51], we adopt both structural and feature-level self-supervised losses. Since the magnitudes of ${\mathcal{L}}_{\text{stru}}$ and ${\mathcal{L}}_{\text{feat}}$ are comparable, we simply aggregate them without additional weighting. This encourages the model to learn robust structural representations by distinguishing between different structural augmentations. The feature self-supervised loss ${\mathcal{L}}_{\text{feat}}$ involves masking portions of the vertex features and reconstructing them based on domain-specific information, similar to the approach used in GraphMAE [52]. This mechanism ensures that the model effectively captures both the structural dependencies and the semantic features of the vertices.
@@ IV-F Discussions
We compare the proposed Hyper-FM with classical graph foundation models, focusing on core design differences and framework positioning, starting with Graph COordinators for PrEtraining (GCOPE)[53].
First, in vertex feature generation: GCOPE uses BoW [16] or TF-IDF [50] for feature extraction, then unifies dimensions via SVD [17], failing to fully leverage advanced language models (LMs) and ignoring inherent data relationships. Hyper-FM, by contrast, adopts LMs to extract fine-grained textual features, and fine-tunes LMs via hierarchical domain representations and domain prediction tasks, integrating structural information into features. Second, in multi-domain association structure generation: GCOPE concatenates multi-domain structures using virtual vertices connected to all domain vertices, causing information confusion among semantically distinct vertices. Hyper-FM constructs hierarchical multi-domain hypergraphs via clustering, strengthening connections between semantically similar vertices and linking domain vertices to clustered ones, enabling orderly cross-domain knowledge transfer. Third, in scalability and transferability: GCOPE’s direct structure concatenation struggles with large-scale hypergraphs, and its domain-specific trainable virtual vertex embeddings require full retraining for new domains. Hyper-FM samples substructures per epoch to handle large-scale data, and generates virtual cluster/bond vertex features via aggregation, avoiding extra trainable parameters and supporting seamless new domain integration.
Beyond GCOPE, we further compare Hyper-FM with LLaGA[31] and GraphGPT[30], which represent an LLM-centered paradigm. LLaGA reorganizes graph nodes into structure-aware sequences and maps them to LLM token space via a projector, adapting graph data to LLM inputs with LLM as the core. GraphGPT integrates LLMs with graph knowledge through graph instruction tuning, including text-graph grounding and dual-stage tuning, enabling LLMs to understand graph structures. Both models rely on LLMs to capture graph information by “tokenizing” graphs. Hyper-FM, however, takes GNNs as the core framework; LMs only serve to align textual features across domains, assisting GNNs in learning better hypergraph knowledge—representing an essentially different design paradigm.
Overall, Hyper-FM outperforms GCOPE in vertex feature generation, multi-domain structure construction, and scalability. It also differs fundamentally from LLaGA/GraphGPT in framework positioning, setting a new benchmark for hypergraph foundation models.